OpenStudy (anonymous):

Need help with 3 questions?

OpenStudy (anonymous):

I'm here to help!

OpenStudy (anonymous):

1+1 = 2 , 2+2 = 4, 3+3 = 6

OpenStudy (chrisplusian):

cosine is adjacent over hypotenuse. The upper case A in problem 8 indicates the problem wants to know the cosine value for angle A. Agree?

OpenStudy (chrisplusian):

You there?

OpenStudy (chrisplusian):

do you just want the answer?

OpenStudy (anonymous):

Got the answers. Nevermind,.

OpenStudy (chrisplusian):

I was just asking cause if you wanted me to help you through them I will. I just didn't know if you wanted to compare your answers or if you were needing help understanding them

OpenStudy (anonymous):

k